OpenStudy (anonymous):

Show all work. SOLVE for P. r= 4Q - 5P -------- 9

OpenStudy (anonymous):

okay. what would you do to isolate P?

OpenStudy (anonymous):

i would multiply both sides by 9?

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

correct! what do you get when you do that?

OpenStudy (anonymous):

9r=4Q-5P

OpenStudy (anonymous):

very good. what would be the next step?

OpenStudy (anonymous):

Subtract 4Q from both sides?

OpenStudy (anonymous):

sure, that works

OpenStudy (anonymous):

9r-4Q=-5P divide by -5?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thats right!

OpenStudy (anonymous):

9r-4Q/-5=P ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

or 4Q-9r/5 =P

OpenStudy (anonymous):

yeah. If you want you can distribute a -1 to the numerator instead of the denominator to make it 'prettier' but it's still right regardless

OpenStudy (anonymous):

yeah, that's correct also

OpenStudy (anonymous):

yup, P = 4Q-9R/5 that is the elegant way to write it finally.

OpenStudy (anonymous):

great work! high five, everyone!

OpenStudy (anonymous):

hey what do you mean by a-1?

OpenStudy (anonymous):

I GIVE 10

OpenStudy (anonymous):

2 high 5's that is

OpenStudy (anonymous):

he meant "a" -1 or, "one" -1

OpenStudy (anonymous):

yes I got it :) see? all you had to do was work it out on your own. good job!

OpenStudy (anonymous):

so how should I write the final answer where it looks nice?

OpenStudy (anonymous):

\[p = \frac{4Q-9R}{5}\]

OpenStudy (anonymous):

I know it is just nice to know that there is somewhere here with me to see if I am doing it right online math class is hard and I am doing my finals and I need to pass this time TO MUCH $$$$$$$$$$$$$$$$$$$$ and I am broke broke broke like the world

OpenStudy (anonymous):

$$$$$$$$4

OpenStudy (anonymous):

yes, i understand.

OpenStudy (anonymous):

cost to much money it did not take the $$$

OpenStudy (anonymous):

Thanks everyone Great Job look out for me I need good answers ok : ) xoxoxox